OpenStudy (anonymous):

I've done everything I possibly could do and I can't figure this out. Can someone please help me with my last problem. Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x^2 + 13x + 15 2x^2 + 27x + 13 5x^2 − 2x − 7 10x^2 − 7x − 3 8x^2 − 11x + 3

OpenStudy (anonymous):

I can't factor trinomials completely.

OpenStudy (anonymous):

for the last (x-1)(x-3/8)

OpenStudy (anonymous):

oh and multimply by 8

OpenStudy (anonymous):

last 3 has factor

OpenStudy (anonymous):

I have to explain.

OpenStudy (callisto):

For the first one, 2x^2 + 13x + 15 = 2x^2 + 3x + 10x + 15 = x(2x+3) + 5(2x+3) = (x+5)(2x+3) For the second one, 2x^2 + 27x + 13 = 2x^2+ x + 26x + 13 = x(2x +1) + 13 (2x+1) = ...? For the third one, 5x^2 − 2x − 7 = 5x^2 + 5x -7x -7 = 5x (x+1) - 7(x+1) = ...? For the fourth one, 10x^2 − 7x − 3 = 10x^2 -10x - 3x -3 = 10x ( x+1) -3(x+1) = ...? For the last one, 8x^2 − 11x + 3 = 8x^2 - 8x - 3x +3 = 8x(x-1) - 3(x-1) = ...? Basically, you expand the middle term, take out the common factor and group the rest of the terms, and repeat this process. You can observe some pattern via the above work

OpenStudy (anonymous):

3 rd 5(x+1)(x-7/5) 4 th 10(x-1)(x+3/10)

OpenStudy (callisto):

You need not take that 5 out for the 3rd one.. same for the 4th one...